Steve's age is seven-tenths of Paul's age. In 4 years Steve's age will be eight-elevenths of Paul's age. How old is Paul now?

Answer by Maths68(1474)

Let
Present age of Steve = s
Present age of Paul = p
Steve's age is seven-tenths of Paul's age
s=7/10 * p
s=7p/10 ........(1)
In 4 years means we have to add 4 in the present ages of both
s+4=P+4
Then
Steve's age will be eight-elevenths of Paul's age
s+4=8/11 * (p+4)
s+4=8(p+4)/11
s+4=(8p+32)/11
11*(s+4)=8p+32
11s+44=8p+32
11s=8p+32-44
11s=8p-12...............(2)
Put the value of s from (1) to (2)
11(7p/10)=8p-12
77p/10=8p-12
77p=10*(8p-12)
77p=80p-120
77p-80p=-120
-3p=-120
Divide by -3 both sides of the equation
-3p/-3=-120/-3
p=40
Put the value of p in (1)
s=7p/10 ........(1)
s=7(40)/10
s=280/10
s=28
Present age of Steve = 28 years old
Present age of Paul = 40 years old